OpenStudy (anonymous):

For the series, determine if its convergent or divergent. if its convergent, find its limit.

OpenStudy (anonymous):

\[\sum_{n=2}^{\infty} \frac{ n }{ \ln(n) }\]

OpenStudy (kewlgeek555):

Whoa! Is this what I'm going to be seeing in college?! -Fixes glasses-

OpenStudy (anonymous):

yeah this is calculus III

OpenStudy (anonymous):

for university.

OpenStudy (anonymous):

Use the limit test: if \[\lim_{n \rightarrow \infty}a_n\neq 0\] then \[\sum a_n\] does not converge

OpenStudy (kewlgeek555):

Calculus. I hope its better than geometry. I hope its like algebra. I love algebra.